How do I find the numerical answer for log3(18) - log3(2)? We are not allowed to use calculators.
Found 2 solutions by scott8148, nerdybill:
Answer by scott8148(6628)

subtracting logs means dividing their arguments

log3(18 / 2) = log3(9) = 2 ___ 3^2 = 9

Answer by nerdybill(7384)